OpenStudy (amistre64):

Recurrsion equations: X{n+1} = (1/2) X{n} + 10*(1/4^n). Any idea how to find the general and particular solutions to this thing?

OpenStudy (anonymous):

just need particular...I've tried over 30 different ways...grrrhhhh!

OpenStudy (anonymous):

it seems that I'm getting close but no cigar, lol!

OpenStudy (amistre64):

Do you recall what X{0} is? so that we can make the table?

OpenStudy (anonymous):

100...I have the table

OpenStudy (anonymous):

60, 32.5, 16.875, 8.59375, 4.3359375, 2.177734375, 1.091308594

OpenStudy (amistre64):

if X{0} = 100 then: X{1} = 50 + 10*1 = 60..... ok. that looks right then :)

OpenStudy (amistre64):

X{0} = 100; X{1} = 60; X{2} = 32.5; X{3} = 16.875; X{4} = 8.59375; X{5} = 4.3359375; X{6} = 2.177734375; X{7} = 1.09130859375; X{8} = 0.5462646484375;

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

and what did you find for a general solution?

OpenStudy (anonymous):

haven't : (

OpenStudy (amistre64):

oh, is particular ahead of general then?

OpenStudy (anonymous):

usually general but this question only asked for particular

OpenStudy (amistre64):

A{n+1} = (1/2) A{n} + 5*(1/2)^(n) is what Wolfram pops out, but it gives no explanation as to how it gets it :)

OpenStudy (anonymous):

ok thanks alot!

OpenStudy (anonymous):

hey man have some owlet than u will have some idea why u ,an intelligent person ask for help

OpenStudy (amistre64):

i dont smoke, but thanx? asking questions is the sign of intelligence, otherwise, we would only know what we know and never expand to greater horizons..

OpenStudy (anonymous):

but u r kinda poor in an IQ as i know.

OpenStudy (amistre64):

:) that may be true

OpenStudy (anonymous):

its almost tru cuz last time u told u left ur iq in los angeels

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

left it? nah....LOST it :)

OpenStudy (anonymous):

the same thing.

OpenStudy (anonymous):

isnt it?

OpenStudy (anonymous):

u lost cuz u left.

OpenStudy (amistre64):

your the one with the IQ ;) you tell me... to leave is to know where to get it from; too lose is to never find again

OpenStudy (anonymous):

so than u r totally dull in an IQ.

OpenStudy (amistre64):

Yes, I am a vaccuum when it comes to IQ.

OpenStudy (amistre64):

in knowing nothing, I leave myself open to know everything; and in knowing everything, I become nothing....

OpenStudy (anonymous):

well u r everything for those guys whom u helped,,but u are nothing for me lol...

OpenStudy (anonymous):

just kidding lol

OpenStudy (anonymous):

you are genius...

OpenStudy (anonymous):

keep it up!the next einstien .

OpenStudy (amistre64):

:)